# Volatility Forecasting Using Similarity-based Parameter Correction and Aggregated Shock Information

### v2

@@ 1 Introduction
In the context of foreign exchange rate market structure, Dominguez and Panthaki [2006] speculate “[w]hether news is scheduled or non-scheduled its influence on exchange rates may be related to the state of the market at the time of the news arrival. News that arrives during periods of high uncertainty may have different effects on the exchange rate, than news that arrives in calmer periods.” The authors also note that non-scheduled news may require more time for markets to digest, leading to greater heterogeneity (including but not limited to higher dispersion) of responses. We take inspiration from these arguments, developing a method suitable to the conditions that typically accompany news shocks.
In this work we focus on the second central moment of a time series, the volatility. One of the most important stochastic phenomena of positively-valued time series $(P_{t})_{t\in\mathbb{N}}$, especially financial time series, is the volatility of the return series $(r_{t})_{t\in\mathbb{N}}$. A financial asset’s price series may exhibit behavior that makes inapplicable and uninterpretable the traditional methods of time series analysis. In contrast, the return series is scale-free [Tsay, 2005], easily-interpreted, and often at least weakly stationary. Our reasons for studying volatility are two-fold. Whereas once returns were thought to react to specific news events, now stock price movements are believed to be overwhelmingly noise-based (see Boudoukh et al. [2019] and references therein), and even if one could construct credible models for describing and forecasting price series and return series, that would not necessarily tell us much about the variability of such forecasts nor enlighten us about the evolution of the variability of $(P_{t})_{t\in\mathbb{N}}$ and $(r_{t})_{t\in\mathbb{N}}$ over time.
@@
The primary methodological tool presented in this work is fixed effect estimation followed by an appropriate procedure for pooling those estimates. The use of fixed effect estimation for the study of structural shocks has a pedigree in macroeconomic analysis (Romer and Romer [1989] cited in Kilian and Lütkepohl [2017]; see also discussion of determinstic exogenous events in Engle and Patton [2001]). We employ fixed effect estimation on the basis of a well-established conceptual assumption that shocks of economic time series can be modeled as mixtures, in particular, mixtures of ordinary innovations and rare events (see Phillips [1996] and references therein). In the forecasting literature, the term “intercept correction” has come to refer to a modeling technique in which nonzero errors are explicitly permitted [Hendry and Clements, 1994, Clements and Hendry, 1998]. They summarize the literature as distinguishing two families of intercept correction: so-called “discretionary” intercept corrections that attempt to account for future events, without hard-coding an ad-hoc adjustment into the model specification, and second, “automated” intercept corrections that attempt to adjust for persistent misspecification using past errors. Guerrón-Quintana and Zhong [2017] use weighted subsets of a scalar series’ own past to correct forecast errors by an additive term. Dendramis et al. [2020] a introduces a similarity-based forecasting procedure for time-varying coefficients of a linear model. Foroni et al. [2022] employ a form of intercept correction in order to adjust forecasts to the COVID-19 shock in the spring of 2020 based on the proportional misses of the same model applied to the Great Recession.
@@
Our procedure is a discretionary procedure for intercept correction that incorporates systematically data internal or external to the time series under study. The correction function, as we shall see, involves an optimization step inspired by the causal inference literature. In particular, in Abadie et al. [2010], the authors build upon previous work in causal inference whereby a treatment effect can be estimated via comparison with a synthetic time series that that represents either the treatment or control unit. The synthetic unit is constructed using a convex combination of the donors. The particular convex combination employed is a function of the distance between the time series under study and the donors. Lin and Eck [2021] adapt these methods for the purpose of prediction. Their one-step-ahead forecasts use distance-based-weighting to pool shock estimates from similar series according to the donor series’ similarity to the series under study. Their approach does not take into account the ARCH effects commonly observed in time series, especially financial times series, leaving unaccounted for the variability that accompanies predictions of a heteroskedastic time series. Outside of Lin and Eck [2021], we know of no prior work that both introduces a parametric specification for nonzero errors and introduces a procedure for weighting appropriately the nonzero errors of similar shocks occurring outside the time series under study. Likewise, we are not familiar with any prior work that attempts to account for anticipated nonzero errors using an explicit parametric adjustment, i.e., what we will call a “correction function”.
The method proposed herein is supported by both numerical evidence as well as a real data example. The numerical evidence takes the form of simulations and analysis of the method’s performance across a considerable grid of varying parameters, allowing us to confirm reasonable hypotheses about the method but also uncover some intriguing nuances. The real data example, on the other hand, demonstrates the usefulness of the method by applying it to the most closely-followed democratic political event on earth, the US Presidential Election. We find that the method would have allowed a researcher to predict the volatility of a financial asset that immediately followed the 2016 US Presidential Election in the United States. Our discussion of this example sheds light on some subtleties of the method as well as its robustness to various choices that must be made to construct and assemble donors and appropriate covariates.
@@ 2 Setting
In order to motivate our procedure, we provide a visual illustration. In Figure 1, we show how the aggregation of estimated excess volatilities from donors in the donor pool works when the correction function is a specially-chosen convex combination of fixed effects from the donor pool. Our method assumes a credible, parsimonious parameterization in which the shock is an affine transformation of several key covariates. The key intuition behind this shock parameterization is that as the strength of the linear signal increases relative to the idiosyncratic error, the GARCH estimation of these effects increases in accuracy. From this, it follows that the aggregated shock estimate increases in accuracy.
Refer to caption ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\sigma}^{2}_{adjusted}=\hat{\sigma}^{2}_{unadjusted}+\hat{\omega}^{*}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\omega}^{*}=\sum^{n+1}_{i=2}\pi_{i}\hat{\omega}^{*}_{i}}$
Figure 1: The simulated time series experiences a volatility shock between trading days 1223 and 1224. We have truncated the series after $T^{*}+1=1224$. The GARCH prediction, in red, fails even to approach the volatility spike at $T^{*}+1$, as do some adjusted predictions, which are orange. In contrast, the GARCH forecast adjusted by $\hat{\omega}^{*}$, a convex combination of the estimated shocks in the donor pool, achieves directional correctness as well as a smaller absolute loss in its prediction.
@@ 2.1 A Primer on GARCH
To that end, let $a_{t}=r_{t}-\mu_{t}$, where $\mu_{t}$ is the mean of the log return series $r_{t}$. We thus derive a mean-zero process $(a_{t})_{t\in\mathbb{N}}$ with the property that $\mathbb{E}[a^{2}_{t}]=\mathrm{Var}[a_{t}]$. Under the assumption of time-invariant volatility, the series $a_{t}^{2}$ should exhibit no autocorrelation at any lag $\ell\geq 1$. This assumption motivates tests for so-called ARCH effects, that is, tests for the clustering of volatility. These tests explore the alternative hypothesis that $\sigma_{t}^{2}$ is not only a time-varying parameter but furthermore a function of past squared residuals of the mean model. In particular, the ARCH($m$) model is an autoregressive model in which $\sigma_{t}^{2}$ is a deterministic function of the past $m$ values of $r_{t}^{2}$. The GARCH($m,s$) framework take this one step further by modeling $\sigma_{t}^{2}$ as a linear combination of the past $m$ values of $r_{t}^{2}$ and well as the past $s$ values of $\sigma_{t}^{2}$. In functional form, a GARCH process (sometimes called a strong GARCH process [Francq and Zakoian, 2019, p. 19]) is given by
@@ 2.2 Model setup
We will suppose that a researcher has multivariate time series data $\textbf{y}_{i,t}=(r_{i,t}$, $\textbf{v}_{i,t}$), $t=1,$ $\ldots,T_{i}$, $i=1,\ldots,n+1$, and $\textbf{v}_{i,t}$ is a vector of covariates such that $\textbf{v}_{i,t}|\mathcal{F}_{i,t-1}$ is deterministic. Suppose that the analyst is interested in forecasting the volatility of $r_{1,t}$, the first time series in the collection, which we will denote the time series under study. We require that each time series $\textbf{y}_{i,t}$ is subject to a news event following $T^{*}_{i}\leq T_{i}+1$ and before witnessing $T^{*}_{i}+1$. We are implicitly leveraging the fact that financial assets are heavily traded during market hours, yet only thinly traded (if traded at all) outside market hours. In contrast, the arrival of market-moving news does not obey any such restrictions. In light of the foregoing, we can denote our collection of GARCH volatility equations of interest using the following notation
@@
Let $I(\cdot)$ be an indicator function. Let $T_{i}$ denote the time length of the time series $i$ for $i=1,\ldots,n+1$, and let $T_{i}^{*}$ denote the largest time index prior to the arrival of the news shock, with $T_{i}^{*}<T_{i}$. Let $\delta,\textbf{v}_{i,t}\in\mathbb{R}^{p}$. For $t=1,\ldots,T_{i}$ and $i=1,\ldots,n+1$, the model $\mathcal{M}_{1}$ is defined as
@@
with error structure
@@
where $D^{return}_{i,t}=I(t\in\{T_{i}^{*}+1,...,T_{i}^{*}+L_{i,return}\})$ and $D^{vol}_{i,t}=I(t\in\{T_{i}^{*}+1,...,T_{i}^{*}+L_{i,vol}\})$ and $L_{i,return},L_{i,vol}$ denote the lengths of the log return and volatility shocks, respectively. Let $\mathcal{M}_{0}$ denote the subclass of $\mathcal{M}_{1}$ models such that $\delta\equiv 0$. Note that $\mathcal{M}_{0}$ assumes that $\omega^{*}_{i}$ have no dependence on the covariates and are i.i.d. with $\mathbb{E}[\omega^{*}_{i}]=\mu_{\omega^{*}}$.
2.3 Volatility Profile of a Time Series
We now introduce prediction via distance-based weighting by constructing a profile of a time series’ volatility, the analogue of a covariate matrix in a synthetic control framework. The volatility profile for a given GARCH-X process is nothing more than a vector of observable covariates that parameterize a shock. A collection of $n$ such vectors yields a matrix of $n$ columns. What distinguishes a volatility profile, principally, is that the time-varying parameters that correspond to the volatility profile are non-zero at only the shock times.
Suppose that for each of the $n$ donors, we have available $p$ distinct covariates in the functional form of the shock. The volatility profile could take the form of the $p\times n$ matrix
$\textbf{V}_{t}=\begin{pmatrix}RV_{2,t}&RV_{3,t}&\cdots&RV_{n+1,t}\\ RV_{2,t-1}&RV_{3,t-1}&\cdots&RV_{n+1,t-1}\\ \vdots&\vdots&\ddots&\vdots\\ IV_{2,t}&IV_{3,t}&\cdots&IV_{n+1,t}\\ IV_{2,t-1}&IV_{3,t-1}&\cdots&IV_{n+1,t-1}\\ \vdots&\vdots&\ddots&\vdots\\ |a_{2,t}|&|a_{3,t}|&\cdots&|a_{n+1,t}|\\ |a_{2,t-1}|&|a_{3,t-1}|&\cdots&|a_{n+1,t-1}|\\ \vdots&\vdots&\ddots&\vdots\\ Volume_{2,t}&Volume_{3,t}&\cdots&Volume_{n+1,t}\\ Volume_{2,t-1}&Volume_{3,t-1}&\cdots&Volume_{n+1,t-1}\\ \vdots&\vdots&\ddots&\vdots\\ \Delta RV_{2,t}&\Delta RV_{3,t}&\cdots&\Delta RV_{n+1,t}\\ \Delta RV_{2,t-1}&\Delta RV_{3,t-1}&\cdots&\Delta RV_{n+1,t-1}\\ \vdots&\vdots&\ddots&\vdots\\ \end{pmatrix},$
where RV and IV denote realized volatility and implied volatility, respectively, $\Delta$ denotes the lag operator between consecutive realizations of a variable, and Volume denotes the quantity of traded shares or other units of an asset. Covariates chosen for inclusion in a given volatility profile may be levels, log differences in levels, percentage changes in levels, or absolute values thereof, among many choices.
As shown, $\textbf{V}_{t}$ displays ‘balance’ in that $p$ covariates exist for each of the $n$ donors. In practice, missing values, corrupted values, or unacceptably extreme or noisy estimates may necessitate some sort of matrix completion, a problem that we do not tackle in this work. We now turn to the next section, where $\textbf{V}_{t}$ is employed in a procedure to arrive at a forecast adjustment.
@@ 3.1 Forecasting
A GARCH model is an ARMA on the squares of the observed scalar time series $a^{2}_{t}$ [Tsay, 2005, Francq and Zakoian, 2019, p. 18, p. 46], assuming that $a_{t}$ satisfies fourth-order stationarity. This fact matters for forecasting because the $h$-step-ahead forecasting function for GARCH model is, just like for an ARMA model, the conditional expectation function, $\mathbb{E}[\sigma^{2}_{i,T^{*}+h}|\mathcal{F}_{T^{*}}]$, or practically speaking, the estimate thereof, $\hat{\mathbb{E}}[\sigma^{2}_{i,T^{*}+h}|\mathcal{F}_{T^{*}}]$ [Zivot, 2009]. Here we have presented one-step-ahead forecasts for a GARCH-X(1,1), without loss of generality. For $h=2,3,4,...$, the conditional expectation is computed recursively, as is standard for iterative autoregressive forecasts.

### v3

@@ 1 Introduction
In the context of foreign exchange rate market structure, Dominguez and Panthaki [2006] speculate “[w]hether news is scheduled or non-scheduled its influence on exchange rates may be related to the state of the market at the time of the news arrival. News that arrives during periods of high uncertainty may have different effects on the exchange rate, than news that arrives in calmer periods.” The authors also note that non-scheduled news may require more time for markets to digest, leading to greater heterogeneity (including but not limited to higher dispersion) of responses. We take inspiration from these arguments, developing a method suitable to the conditions that typically accompany news shocks. Central to our endeavor is the observation that news shocks have both a qualitative (the description of the event as well as future contingents surrounding it) and a statistical incarnation, i.e. the reaction of quantitative phenomena in response to the qualitative information.
In this work we focus on the second central moment of a time series, the volatility. One of the most important stochastic phenomena of positively-valued time series $(P_{t})_{t\in\mathbb{N}}$, especially financial time series, is the volatility of the return series $(r_{t})_{t\in\mathbb{N}}$. A financial asset’s price series may exhibit behavior that makes inapplicable and uninterpretable the traditional methods of time series analysis. In contrast, the return series is scale-free [Tsay, 2005], easily-interpreted, and often at least weakly stationary. Our reasons for studying volatility are two-fold. Whereas once returns were thought to react to specific news events, now stock price movements are believed to be overwhelmingly noise-based (see Boudoukh et al. [2019] and references therein), and even if one could construct credible models for describing and forecasting price series and return series, that would not necessarily tell us much about the variability of such forecasts nor enlighten us about the evolution of the variability of $(P_{t})_{t\in\mathbb{N}}$ and $(r_{t})_{t\in\mathbb{N}}$ or under rapidly changing conditions.
@@
The primary methodological tool presented in this work is fixed effect estimation followed by a distance-based weighting procedure for pooling those estimates. The use of fixed effect estimation for the study of structural shocks has a pedigree in macroeconomic analysis (Romer and Romer [1989] cited in Kilian and Lütkepohl [2017]; see also discussion of determinstic exogenous events in Engle and Patton [2001]). We employ fixed effect estimation on the basis of a well-established conceptual assumption that shocks of economic time series can be modeled as mixtures, in particular, mixtures of ordinary innovations and shocks due to rare events (see Phillips [1996] and references therein). In the forecasting literature, the term “intercept correction” has come to refer to a modeling technique in which nonzero errors are explicitly permitted [Hendry and Clements, 1994, Clements and Hendry, 1998]. They summarize the literature as distinguishing two families of intercept correction: so-called “discretionary” intercept corrections that attempt to account for future events, without hard-coding an ad-hoc adjustment into the model specification, and second, “automated” intercept corrections that attempt to adjust for persistent misspecification using past errors. Guerrón-Quintana and Zhong [2017] use weighted subsets of a scalar series’ own past to correct forecast errors by an additive term. Dendramis et al. [2020] a introduces a similarity-based forecasting procedure for time-varying coefficients of a linear model. Foroni et al. [2022] employ a form of intercept correction in order to adjust forecasts to the COVID-19 shock in the spring of 2020 based on the proportional misses of the same model applied to the Great Recession.
@@
Our procedure is a discretionary procedure for intercept correction that incorporates systematically data internal or external to the time series under study. The correction function, as we shall see, includes a convex combination of fixed effect estimates from a donor pool. In the causal inference literature, a donor pool is a set of units of observation that share distributional properties with the prime object of analysis. In Abadie et al. [2010], the authors build upon previous work in causal inference whereby a treatment effect can be estimated via comparison with a synthetic time series that represents the control. In their setting, the synthetic unit is constructed using a convex combination of the donors. The particular convex combination employed is a function of the distance between the time series under study and the donors. Donors that are closer to the time series under study are, ceteris paribus, accorded greater weight. Conversely, donors that are unlike the time series under study receive little to no weight. Lin and Eck [2021] adapt the notion of a donor pool as well as distance-based weighting for the purpose of prediction. Their one-step-ahead forecasts use distance-based-weighting to aggregate shock estimates from the donor pool’s time series according to the similarility to the time series under study. Their approach does not take into account the ARCH effects commonly observed in time series, especially financial times series, leaving unaccounted for the variability in the variability of time series forecasts. Outside of Lin and Eck [2021], we know of no prior work that both introduces a parametric specification for nonzero errors and introduces a procedure for weighting appropriately the nonzero errors of similar shocks occurring outside the time series under study. Likewise, we are not familiar with any prior work that attempts to account for anticipated nonzero errors using an explicit parametric adjustment, i.e., what we will call a “correction function”.
The method proposed herein is supported by both numerical evidence as well as a real data example. The numerical evidence takes the form of simulations and analysis of the method’s performance across a considerable grid of varying parameters, allowing us to evaluate reasonable hypotheses about the method but also uncover some intriguing nuances. The real data example, on the other hand, demonstrates the usefulness of the method by applying it to the most closely-followed democratic political event on earth, the US Presidential Election. We find that the method would have allowed a researcher to predict the volatility of a financial asset that immediately followed the 2016 US Presidential Election in the United States. Our discussion of this example sheds light on some subtleties of the method as well as its robustness to various choices that must be made to construct and assemble donors and appropriate covariates.
@@ 2 Setting
In order to motivate our procedure, we provide a visual illustration. In Figure 1, we show how the aggregation of estimated shock-induced excess volatilities from donors in the donor pool works when the correction function is a specially-chosen convex combination of fixed effects from the donor pool. Our method assumes a credible, parsimonious parameterization in which the shock is an affine transformation of several key covariates. The key intuition behind this shock parameterization is that as the strength of the linear signal increases relative to the idiosyncratic error, the GARCH estimation of these effects increases in accuracy. From this, it follows that the aggregated shock estimate increases in accuracy.
Figure 1: The simulated time series experiences a volatility shock between trading days 1,656 and 1,657. The GARCH prediction, in red, fails even to approach the volatility spike at $T^{*}+1$, as do several adjusted predictions, which are orange. In contrast, the GARCH forecast adjusted by $\hat{\omega}^{*}=\sum^{n+1}_{i=2}\pi_{i}\hat{\omega}^{*}_{i}$, a convex combination of the estimated shocks in the donor pool, achieves directional correctness as well as a smaller absolute loss in its prediction. The pink vertical line serves to indicate the adjustment of size $\hat{\omega}^{*}$ that allows the blue bullseye to more closely approach the ground truth.
@@ 2.1 A Primer on GARCH
To that end, let $a_{t}=r_{t}-\mu_{t}$, where $\mu_{t}$ is the mean of the log return series $r_{t}$. We thus derive a mean-zero process $(a_{t})_{t\in\mathbb{N}}$ with the property that $\mathbb{E}[a^{2}_{t}]=\mathrm{Var}[a_{t}]$. Under the assumption of time-invariant volatility, the series $a_{t}^{2}$ should exhibit no autocorrelation at any lag $\ell\geq 1$. This assumption motivates tests for ARCH effects, that is, tests for the clustering of volatility. These tests explore the alternative hypothesis that $\sigma_{t}^{2}$ is not only a time-varying parameter but furthermore a function of past squared residuals of the mean model. In particular, the ARCH($m$) model is an autoregressive model in which $\sigma_{t}^{2}$ is a deterministic function of the past $m$ values of $r_{t}^{2}$. The GARCH($m,s$) framework take this one step further by modeling $\sigma_{t}^{2}$ as a linear combination of the past $m$ values of $r_{t}^{2}$ and well as the past $s$ values of $\sigma_{t}^{2}$. In functional form, a GARCH process (sometimes called a strong GARCH process [Francq and Zakoian, 2019, p. 19]) is given by
@@ 2.2 Model setup
We will suppose that a researcher has multivariate time series data $\textbf{y}_{i,t}=(r_{i,t}$, $\textbf{v}_{i,t}$), $t=1,$ $\ldots,T_{i}$, $i=1,\ldots,n+1$, and $\textbf{v}_{i,t}$ is a vector of covariates such that $\textbf{v}_{i,t}|\mathcal{F}_{i,t-1}$ is deterministic. Suppose that the analyst is interested in forecasting the volatility of $r_{1,t}$, the first time series in the collection, which we will denote the time series under study. We require that each time series $\textbf{y}_{i,t}$ is subject to an observed news event following $T^{*}_{i}\leq T_{i}+1$ and before witnessing $T^{*}_{i}+1$. We are implicitly leveraging the fact that financial assets are heavily traded during market hours, yet only thinly traded (if traded at all) outside market hours. In contrast, the arrival of market-moving news does not obey any such restrictions. In light of the foregoing, we can denote our collection of GARCH-X volatility equations of interest using the following notation
@@
Let $I(\cdot)$ be an indicator function. Let $T_{i}$ denote the time length of the time series $i$ for $i=1,\ldots,n+1$, and let $T_{i}^{*}$ denote the largest time index prior to the arrival of the news shock, with $T_{i}^{*}<T_{i}$, to ensure that there is at least one post-shock realization for each series $i$. Let $\delta,\textbf{v}_{i,t}\in\mathbb{R}^{p}$. Let $\mathcal{F}$ with a single-variable subscript denote a univariate, time-invariant $\sigma$-algebra, and $\mathcal{F}_{i,t}$ denote the canonical product filtration for donor $i$ at time $t$. Let $D^{return}_{i,t}=I(t\in\{T_{i}^{*}+1,...,T_{i}^{*}+L_{i,return}\})$ and $D^{vol}_{i,t}=I(t\in\{T_{i}^{*}+1,...,T_{i}^{*}+L_{i,vol}\})$ and $L_{i,return},L_{i,vol}$ denote the lengths of the log return and volatility shocks, respectively. For $t=1,\ldots,T_{i}$ and $i=1,\ldots,n+1$, the model $\mathcal{M}_{1}$ is defined as
@@
with time-invariant error structure
@@
Let $\mathcal{M}_{0}$ denote the subclass of $\mathcal{M}_{1}$ models such that $\delta\equiv 0$. Note that $\mathcal{M}_{0}$ assumes that $\omega^{*}_{i}$ have no dependence on the covariates and are i.i.d. with $\mathbb{E}[\omega^{*}_{i}]=\mu_{\omega^{*}}$, where the lack of an index $i$ on $\mu_{\omega^{*}}$ indicates that it is shared across donors.
We now introduce how a researcher can carry out predictions via distance-based weighting. The first step is the gathering of covariates that will represent the $p$-dimensional space in which weights are generated. The set of covariates for our models are nothing than a vector of observable covariates that parameterize a shock. A collection of $n$ such vectors yields a matrix of $n$ columns, $\textbf{V}_{t}$. Formally speaking, the shocks are given by the equation
$\vec{\omega}=\vec{\mu}_{\omega^{*}}+\textbf{V}_{t}\delta+\vec{u}_{t}$
where were have suppressed the $D^{vol}_{i,t}$ notation for simplicity. Suppose that for each of the $n$ donors, we have available $p$ distinct covariates in the functional form of the shock. The covariate set could take the form of a $p\times n$ matrix including realized volatility and implied volatility, as well the volume of traded assets on any day preceding the shock. Covariates chosen for inclusion in a given volatility profile may be levels, log differences in levels, percentage changes in levels, or absolute values thereof, among many choices.
Ideally, $\textbf{V}_{t}$ will display ‘balance’ in that $p$ covariates exist for each of the $n$ donors. In practice, missing values, corrupted values, or unacceptably extreme or noisy estimates may necessitate some sort of matrix completion, a problem that we do not tackle in this work. We now turn to the next section, where $\textbf{V}_{t}$ is employed in a procedure to arrive at a forecast adjustment.
@@ 3.1 Forecasting
A GARCH model is an ARMA on the squares of the scalar time series $a^{2}_{t}$ [Tsay, 2005, Francq and Zakoian, 2019, p. 18, p. 46], assuming that $a_{t}$ satisfies fourth-order stationarity. This fact matters for forecasting because the $h$-step-ahead forecasting function for a GARCH model is, just like for an ARMA model, the conditional expectation function, $\mathbb{E}[\sigma^{2}_{i,T^{*}+h}|\mathcal{F}_{T^{*}}]$, or practically speaking, the estimate thereof, $\hat{\mathbb{E}}[\sigma^{2}_{i,T^{*}+h}|\mathcal{F}_{T^{*}}]$ [Zivot, 2009]. Here we have presented one-step-ahead forecasts for a GARCH-X(1,1), without loss of generality. For $h=2,3,4,...$, the conditional expectation is computed recursively, as is standard for iterative autoregressive forecasts.